Circular No. 8072

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003aq AND 2003ar Further to IAUC 8064, T. Boles reports his discovery of two apparent

supernovae on unfiltered CCD images:

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003aq Feb. 6.070 14^{h}10^{m}07.^{s}24 +17^{o}37′05′′.3 18.5 4′′.6 E, 8′′.1 N
2003ar Feb. 6.162 16 02 08.11 +70 24 55.5 17.9 4′′.8 E, 5′′.9 S

Further magnitudes from Boles’ images: SN 2003aq in NGC 5490C, 2002 Apr. 6 UT, [19.0; May 30, [19.5; 2003 Feb. 14.119, 18.0. SN 2003ar in MCG +12-15-47, 2002 Sept. 7, [19.0; 2003 Jan. 23, [19.5; Feb. 14.142, 16.8. SN 2003aq and SN 2003ar are both absent from second Palomar Sky Survey red and blue plates.

2003 CP_{20}

MPEC 2003-C63 contains observations, orbital elements (a = 0.76 AU, e = 0.29, i = 25^{o}, H = 16.3), and an ephemeris for an object, discovered near its maximum possible elongation (76^{o}) on Feb. 11 by LINEAR, that is the first minor planet with a confirmed aphelion distance < 1 AU (Q = 0.978 AU). The minimum possible distance from the earth is currently 0.19 AU (but passages within 0.05 AU of Venus are occurring).

IGR J16318–4848 N. Schartel, M. Ehle, M. Breitfellner, M. Guainazzi, P. Rodriguez-

Pascual, M. Santos-Lleo, P. Calderon-Riano, P. Munuera-Gallardo, B. Gonzalez-Garcia, and R. Perez-Martinez, XMM-Newton Science Operations Centre, report that an XMM-Newton target-of-opportunity observation on Feb. 10.7 UT shows a bright source at α = 16^{h}31^{m}48^{s}.6, δ = −48^{o}49′00′′ (equinox 2000.0; uncertainty ∼ 4′′) with both EPIC cameras (p-n and MOS) during a 25000-s exposure. This position is consistent with that measured for IGR J16318–4848 by INTEGRAL (IAUC 8063). The p-n count rate shows clear evidence for variability covering a range between ∼ 0.2 and 0.75 counts/s. Images from both EPIC cameras, together with a light curve and spectrum from the p-n camera, are provided at http://xmm.vilspa.esa.

es/external/xmm news/items/IGR/index.shtml.

2003 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green